# Weaving Geodesics and New Phenomena in Horocyclic Dynamics

### v2

@@ Abstract.
The study of horospherical flows on hyperbolic manifolds dates back to Hedlund in the 1930s [Hed36] and has played an important role in the development of modern homogeneous dynamics. In the finite-volume (and geometrically finite) cases, both the measure-theoretic and topological properties of the flow have been extensively studied, revealing a remarkable degree of rigidity (see, e.g. [Fur73, DS84, Bur90, Rob03, Rat91]). In contrast, the behavior of horospherical flows in the general geometrically infinite setting remains much less well understood.
Until recently the only explicitly described examples of horocyclic orbit closures on orientable hyperbolic surfaces were “trivial” — either the full non-wandering set for the horocyclic flow or single closed horocycles. While the existence of other, more intricate, orbit closures was known for decades, none have been described in detail, leaving much mystery as to their potential regularity and rigidity properties (c.f. [DS00, CM10, GL17, Mat16, Bel18b, Led97, Led98]).
In recent works [FLM23, FLM24], an explicit description of all horocyclic orbit closures was given in the setting of $\mathbb{Z}$-covers of compact hyperbolic surfaces. These orbit closures were shown to be highly irregular; their structural features depend in a delicate way on the geometry of the underlying surface. Intriguingly, all had integer Hausdorff dimension. They are also non-minimal.
@@
As a consequence of our constructions, we provide the first counterexamples to the horospherical infinite measure rigidity phenomenon, which has been observed in a vast variety of settings, where every horospherically invariant ergodic Radon measure is either quasi-invariant under the geodesic flow (à la Babillot–Ledrappier) or supported on a single closed orbit; see [Bur90, Rob03, Sar04, LS07, Sar10, OP17, Sar19, LL22, Lan21, LLLO23] and Remark 2.8.
@@
Recall that a non-empty $N$-invariant closed set $F\subseteq\Sigma$ is called $N$-minimal if all $N$-orbits in $F$ are dense in $F$. A characterization of points with dense horocyclic orbits in $\mathcal{E}$ is given by [Ebe77, Dal00] where it was shown that $\overline{Nx}\neq\mathcal{E}$ if and only if the geodesic ray $A_{+}x$ is quasi-minimizing, that is, $d_{\mathsf{T}^{1}\Sigma}(a_{t}x,x)\geq t-c$ for some $c\geq 0$ and every $t\geq 0$. As a consequence, $\mathcal{E}$ is $N$-minimal if and only if $\Gamma$ is convex co-compact.
Studying the different possible trajectories of quasi-minimizing rays and their “efficiency” has turned out to be key in the analysis of horocyclic orbit closures. Drawing on techniques developed in [FLM23, FLM24] and inspired by examples introduced by Alexandre Bellis in [Bel18a, §1.5.1], we provide a recipe for tailoring geometrically infinite surfaces supporting horocycles with prescribed recurrence properties. Our main results are the following:
@@
  • Our surfaces are extremely sparse; the injectivity radius along all diverging geodesic rays tends to infinity. This implies, in particular, that the tameness conditions imposed in [Sar10] to deduce measure rigidity cannot be removed. Equivalent geometric conditions appear in [LL22].
@@
While we rely on techniques developed in [FLM23, FLM24], we will only make use of several elementary insights and lemmas from said papers. Our proof is fairly self-contained and requires no prior knowledge or understanding of the results in the $\mathbb{Z}$-cover setting.
@@ 1.1. Loom Surfaces
Note that under the above conditions the function $\text{Inj--rad}:\Sigma_{s}\to(0,\infty)$, assigning the injectivity radius at a point, is a proper map. In particular, injectivity radius tends to infinity along any diverging geodesic ray.
@@ 1.2. Tight map and stretch lamination
As the composition of $1$-Lipschitz maps, $\tau$ is itself $1$-Lipschitz. Let $\lambda=p(Ax_{0})\cup p(Ax_{1})$ and $\mathsf{T}_{+}^{1}\lambda=Ax_{1}\cup Ax_{2}$. Observe that $\tau$ is strictly contracting away from $\lambda$ and is isometric along each of its components. In particular, $p(Ax_{0})$ and $p(Ax_{1})$ are isometrically embedded in $\Sigma$, and $Ax_{0}$ and $Ax_{1}$ are isometrically embedded in $\mathsf{T}^{1}\Sigma$.
@@ 1.3. Slack
We consider the notion of $\tau$-slack introduced in [FLM24, §3] which measures how “efficiently” a path progresses towards the end at $\infty$:
@@ 1.4. Busemann-type Function
$\beta(y)=\tau(y)-\mathscr{S}_{+}(A_{+}y)=\lim_{t\to+\infty}\tau(a_{t}y)-t.$
It is upper semi-continuous, as a decreasing limit of continuous functions, and also $N$-invariant, see [FLM23, Lemma 6.1]. Therefore, for all $y\in\mathsf{T}^{1}\Sigma_{s}$ we have
@@ 1.5. Weaving Lemma
See the first part of the proof of [FLM23, Thm. 3.4] for more details.
@@
Given $k\in\mathbb{N}$ we denote by $\eta_{k}^{+}$ the unique geodesic in $\Sigma_{s}$ which is backward asymptotic to $A_{-}x_{0}$ and forward asymptotic to $A_{+}x_{1}$ crossing once from $J_{s}\times\{0\}$ to $J_{s}\times\{1\}$ through $\partial D_{h_{k}}(s_{k})$. We call $\eta_{k}^{+}$ a crossing. We similarly have $\eta_{k}^{-}$, the symmetric geodesic passing from $J_{s}\times\{1\}$ to $J_{s}\times\{0\}$ through $\partial D_{h_{k}}(s_{k})$, see Figure 3.
Figure 3.
Notice that by the symmetry of the construction we have $\eta_{k}^{\pm}\cap\partial D_{h_{k}}(s_{k})=\{s_{k}+ih_{k}\}$.
@@
A weaving pattern is a any subset of $\mathbb{N}$ thought of as a strictly increasing sequence of indices $W=\{k_{1}<k_{2}<$...\}.
A weaving geodesic ray with weaving pattern $W$ is an infinite geodesic ray $A_{+}w$ given by pulling tight the following concatenation:
$\alpha_{0}\ast\eta^{\sigma_{1}}_{k_{1}}\ast\eta^{\sigma_{2}}_{k_{2}}\ast\eta^{\sigma_{3}}_{k_{3}}$...
where:
  1. (1)
    $\sigma_{k}\in\{+,-\}$ are alternating signs;
  2. (2)
    $\alpha_{0}$ is the shortest segment connecting $p(w)$ to $\eta^{\sigma_{1}}_{k_{1}}$;
  3. (3)
    the $\ast$ notation is interpreted as concatenation of subsegments beginning and ending at the unique intersection points of the paths;
  4. (4)
    the weaving pattern is allowed to be finite in which case we append, at the end, an infinite subray of either $A_{+}x_{0}$ or $A_{+}x_{1}$, as needed.
$\eta_{W,+}=\widehat{\eta_{k_{1}}^{+}\ast\eta_{k_{2}}^{-}\ast\eta_{k_{3}}^{+}\ast...}$,
where the alternating signs begin with a ++. We similarly define $\eta_{W,-}$. We will refer to such $\eta_{W}^{\pm}$ as weaving geodesics with weaving pattern $W$.
Figure 4. Intersection point of two crossings.
@@
In a loom surface $\Sigma_{s}$, for any $\rho\geq 0$ there exists $S>0$ such that any geodesic ray $A_{+}y$ with slack $\mathscr{S}_{+}(A_{+}y)\leq\rho$ and beginning at $\tau(y)>S$ is weaving.
@@
First note that since both $J_{s}\times\{j\}$ are convex and simply connected, any geodesic ray passing from $J_{s}\times\{j\}$ through $\partial D_{h_{k}}(s_{k})$ to $J_{j+1}$ will either stay indefinitely in the $j+1\mod 2$-side or will have its first return to $J_{s}\times\{j\}$ happen through a different component of $\partial J_{s}$. Therefore, by the uniqueness of geodesic representatives in each homotopy class, it’s enough to show that the sequence of boundary components of $J_{s}$ which the geodesic crosses has increasing indices.
Now for any $\rho\geq 0$ there exists $k_{0}\in\mathbb{N}$ such
@@ 1.6. Slack of a Weaving Geodesic
We invoke the following lemma from [FLM24, Lemma 3.5]:
@@
A fundamental observation, highlighting the utility of the notion of slack, is the following lemma proven in [FLM24, Lemma 3.3]:
@@ 2. Minimal Orbit Closures
$\sum_{j=1}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})=\sum_{j=1}^{\infty}2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{k}))<\infty.$
@@
where $\mathcal{E}$ denotes the non-wandering set for the horocycle flow on $\mathsf{T}^{1}\Sigma_{s}$.
@@
$\liminf_{t\to+\infty}d(a_{t}y,a_{t}w)=0.$
We will make use of the following fact, proven in [FLM23, Cor. 8.3]:
@@
Our strategy for proving Theorem 2.2 will be to show that all the points in $\overline{Nx_{0}}$ are $A$-proximal to $x_{0}$. Two key observations under these assumptions:
Lemma 2.4.
If $\beta(y)=b>$-$\infty$ then $y$ is $A$-proximal to $a_{b}x_{0}$.
@@
(2.1) $\lim_{k\to\infty}d\left(a_{s_{k}}x_{0},a_{s_{k}}x_{1}\right)=0.$
In particular, for all $k$ large enough we have
(2.2) $\mathsf{T}^{1}_{+}\lambda\cap\tau^{-1}\left((s_{k}-\varepsilon,s_{k}+\varepsilon)\right)\subset B_{3\varepsilon}(a_{s_{k}}x_{0}).$
Assume $y$ as in the statement. For any $\varepsilon>0$ there exists $T>0$ large enough such that $a_{t}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$ for all $t\geq T$ (1.4). By the definition of $\beta$ we have $|\tau(a_{t-b}y)-t|\to 0$ as $t\to\infty$. Hence for all $k$ large enough we have that $|\tau(a_{s_{k}-b}y)-\tau(a_{s_{k}}x_{0})|<\varepsilon$ and $a_{s_{k}-b}y$ is in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$. By (2.2) we conclude
$\liminf_{k\to\infty}d(a_{s_{k}-b}y,a_{s_{k}-b}a_{b}x_{0})\leq 3\varepsilon,$
implying the claim. ∎
The second observation is:
@@
For any $\varepsilon>0$ there exists $k_{0}\in\mathbb{N}$ such that any weaving pattern $W$ satisfying $W\geq k_{0}$ will have $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon$.
@@
Recall that the past of $\eta_{W,\pm}$ is asymptotic to $\mathsf{T}^{1}_{+}\lambda$ and that this geodesic has arbitrarily long segments which are arbitrarily close to $\mathsf{T}^{1}_{+}\lambda$ (close to intersection points of the crossings). This is also true for the “untightened” path $\alpha_{W}=\eta_{k_{1}}^{\pm}\ast\eta_{k_{2}}^{\mp}\ast$.... Hence we can compare slacks along exhausting subsegments of $\eta_{W,\pm}$ and $\alpha_{W}$ which have $\tau$-value arbitrarily close at their respective endpoints. The main difference is that the path along $\eta_{W,\pm}$ is shorter. Therefore we have $\mathscr{S}_{+}(\eta_{W,\pm})<\mathscr{S}_{+}(\alpha_{W})$. The inequality
$\mathscr{S}_{+}(\alpha_{W})<\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})$
follows from the additivity and non-negativity of slack. ∎
@@
First, by (2.1) we note that $x_{0}$ and $x_{1}$ are $A$-proximal. Therefore by 2.3 we have $x_{1}\in F=\overline{Nx_{0}}$. Since clearly $x_{1}\notin Nx_{0}$ (the corresponding rays $A_{+}x_{j}$ are not asymptotic) we deduce that $F\neq Nx_{0}$. The strict inclusion $F\subsetneq\mathcal{E}$ follows from $x_{0}$ being quasi-minimizing.

### v3

@@ Abstract.
The study of horospherical flows on hyperbolic manifolds dates back to Hedlund in the 1930s [14] and has played an important role in the development of modern homogeneous dynamics. In the finite-volume (and geometrically finite) cases, both the measure-theoretic and topological properties of the flow have been extensively studied, revealing a remarkable degree of rigidity (see, e.g. [12, 8, 4, 30, 28]). In contrast, the behavior of horospherical flows in the general geometrically infinite setting remains much less well understood.
Until recently the only explicitly described examples of horocyclic orbit closures on orientable hyperbolic surfaces were “trivial” — either the full non-wandering set for the horocyclic flow or single closed horocycles. While the existence of other, more intricate, orbit closures was known for decades, none have been described in detail, leaving much mystery as to their potential regularity and rigidity properties (c.f. [6, 5, 13, 24, 3, 22, 23]).
In recent works [10, 11], an explicit description of all horocyclic orbit closures was given in the setting of $\mathbb{Z}$-covers of compact hyperbolic surfaces. These orbit closures were shown to be highly irregular; their structural features depend in a delicate way on the geometry of the underlying surface. Intriguingly, all had integer Hausdorff dimension. They are also non-minimal.
@@
As a consequence of our constructions, we provide the first counterexamples to the horospherical infinite measure rigidity phenomenon, which has been observed in a vast variety of settings, where every horospherically invariant ergodic Radon measure is either quasi-invariant under the geodesic flow (à la Babillot–Ledrappier) or supported on a single closed orbit; see [4, 30, 32, 21, 33, 25, 34, 19, 20, 18] and Remark 2.9.
@@
Recall that a non-empty $N$-invariant closed set $F\subseteq\Sigma$ is called $N$-minimal if all $N$-orbits in $F$ are dense in $F$. A characterization of points with dense horocyclic orbits in $\mathcal{E}$ is given by [9, 7] where it was shown that $\overline{Nx}\neq\mathcal{E}$ if and only if the geodesic ray $A_{+}x$ is quasi-minimizing, that is, $d_{\mathsf{T}^{1}\Sigma}(a_{t}x,x)\geq t-c$ for some $c\geq 0$ and every $t\geq 0$. As a consequence, $\mathcal{E}$ is $N$-minimal if and only if $\Gamma$ is convex co-compact.
Studying the different possible trajectories of quasi-minimizing rays and their “efficiency” has turned out to be key in the analysis of horocyclic orbit closures. Drawing on techniques developed in [10, 11] and inspired by examples introduced by Alexandre Bellis in [2, §1.5.1], we provide a recipe for tailoring geometrically infinite surfaces supporting horocycles with prescribed recurrence properties. Our main results are the following:
@@
  • Our surfaces are extremely sparse; the injectivity radius along all diverging geodesic rays tends to infinity. This implies, in particular, that the tameness conditions imposed in [33] to deduce measure rigidity cannot be removed. Equivalent geometric conditions appear in [19].
@@
While we rely on techniques developed in [10, 11], we will only make use of several elementary insights and lemmas from said papers. Our proof is fairly self-contained and requires no prior knowledge or understanding of the results in the $\mathbb{Z}$-cover setting.
@@ 1.1. Loom Surfaces
Note that for any sequence $0<h_{k}$ uniformly bounded away from $\frac{\pi}{2}$ there exist $s_{k}\to\infty$, sufficiently large and spread apart, for which $s=(s_{k},h_{k})_{k\in\mathbb{N}}$ defines a loom surface.
Under the above conditions the function $\text{Inj--rad}:\Sigma_{s}\to(0,\infty)$, assigning the injectivity radius at a point, is a proper map. In particular, injectivity radius tends to infinity along any diverging geodesic ray.
@@ 1.2. Tight map and stretch lamination
As the composition of $1$-Lipschitz maps, $\tau$ is itself $1$-Lipschitz. Let $\lambda=p(Ax_{0})\cup p(Ax_{1})$ and $\mathsf{T}_{+}^{1}\lambda=Ax_{0}\cup Ax_{1}$. Observe that $\tau$ is strictly contracting away from $\lambda$ and is isometric along each of its components. In particular, $p(Ax_{0})$ and $p(Ax_{1})$ are isometrically embedded in $\Sigma$, and $Ax_{0}$ and $Ax_{1}$ are isometrically embedded in $\mathsf{T}^{1}\Sigma$.
@@ 1.3. Slack
We consider the notion of $\tau$-slack introduced in [11, §3] which measures how “efficiently” a path progresses towards the end at $\infty$:
@@ 1.4. Busemann-type Function
(1.2) $\beta(y)=\tau(y)-\mathscr{S}_{+}(A_{+}y)=\lim_{t\to+\infty}\tau(a_{t}y)-t.$
It is upper semi-continuous, as a decreasing limit of continuous functions, and also $N$-invariant, see [10, Lemma 6.1]. Therefore, for all $y\in\mathsf{T}^{1}\Sigma_{s}$ we have
@@ 1.5. Weaving Lemma
See the first part of the proof of [10, Thm. 3.4] for more details.
@@
A geodesic ray $A_{+}w$ is said to be weaving with pattern $W=\{k_{1}<k_{2}<$$\subseteq\mathbb{N}$ if $t\mapsto\tau(a_{t}w)$ is monotonically increasing and unbounded, for $t\geq 0$, and
$W=\left\{k\in\mathbb{N}:p(A_{+}w)\cap\partial D_{h_{k}}(s_{k})\neq\emptyset\right\}.$
@@
In a loom surface $\Sigma_{s}$, weaving geodesics are uniquely determined by their weaving pattern and initial point.
In addition, for any $\rho\geq 0$ there exists $S>0$ such that any geodesic ray $A_{+}y$ with slack $\mathscr{S}_{+}(A_{+}y)\leq\rho$ and beginning at $\tau(y)>S$ is weaving.
@@
First note that since both $J_{s}$ copies are convex and simply connected, any geodesic ray passing from $J_{s}\times\{j\}$ through $\partial D_{h_{k}}(s_{k})$ to $J_{s}\times\{j+1\!\!\mod 2\}$ will either stay indefinitely in that side or have its first return to $J_{s}\times\{j\}$ happen through a different component of $\partial J_{s}$. Whenever a weaving geodesic has an infinite weaving pattern, monotonicity of $\tau$-values ensures the indices in the pattern appear in strictly increasing order. Uniqueness of geodesic representatives in each homotopy class relative endpoints implies such a weaving geodesic is uniquely determined.
On the other hand, any geodesic ray entirely contained within one $J_{s}\times\{j\}$ and having unbounded $\tau$-values is necessarily asymptotic to $A_{+}x_{j}$. We may thus deduce the claim for weaving geodesics with a finite pattern as well.
For the second claim, given any $\rho\geq 0$ there exists $k_{0}\in\mathbb{N}$ such
@@
Given $k\in\mathbb{N}$ we denote by $\eta_{k}^{+}$ the unique geodesic in $\Sigma_{s}$ which is backward asymptotic to $A_{-}x_{0}$ and forward asymptotic to $A_{+}x_{1}$ crossing once from $J_{s}\times\{0\}$ to $J_{s}\times\{1\}$ through $\partial D_{h_{k}}(s_{k})$. We call $\eta_{k}^{+}$ a crossing. We similarly have $\eta_{k}^{-}$, the symmetric geodesic passing from $J_{s}\times\{1\}$ to $J_{s}\times\{0\}$ through $\partial D_{h_{k}}(s_{k})$, see Figure 3. Symmetry of the construction implies $\eta_{k}^{\pm}\cap\partial D_{h_{k}}(s_{k})=\{s_{k}+ih_{k}\}$.
Figure 3.
(1.4) $\eta_{W,+}=\widehat{\eta_{k_{1}}^{+}\ast\eta_{k_{2}}^{-}\ast\eta_{k_{3}}^{+}\ast...}$,
where the $\ast$ notation is interpreted as concatenation of subsegments beginning and ending at the unique intersection points of the paths, see Figure 4. We similarly define $\eta_{W,-}$ where the alternating signs of the crossing begin with a ‘-’ sign. Whenever the weaving pattern is finite the concatenation ends with a full subray of the final crossing.
Figure 4. Intersection point of two crossings.
The weaving lemma allows us to draw the following realization lemma:
Lemma 1.7.
Every weaving geodesic $A_{+}w$ with pattern $W$ is given by pulling tight $\alpha_{0}\ast\eta_{W,\sigma}$ where $\alpha_{0}$ is any path connecting $p(w)$ to $\eta^{\sigma_{1}}_{k_{1}}$ fully contained within one $J_{s}\times\{j\}$ and $\sigma\in\{\pm\}$ is chosen appropriately.
@@ 1.6. Slack of a Weaving Geodesic
We invoke the following lemma from [11, Lemma 3.5]:
@@
A fundamental observation, highlighting the utility of the notion of slack, is the following lemma proven in [11, Lemma 3.3]:
@@ 2. Minimal Orbit Closures
$\sum_{j=1}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})=\sum_{j=1}^{\infty}2\ln\cosh(d_{\mathbb{H}^{2}}(0,ih_{j}))<\infty.$
@@
where $\mathcal{E}$ denotes the non-wandering set for the horocyclic flow on $\mathsf{T}^{1}\Sigma_{s}$.
@@
$\liminf_{t\to+\infty}d_{\mathsf{T}^{1}\Sigma_{s}}(a_{t}y,a_{t}w)=0.$
We will make use of the following fact, proven in [10, Cor. 8.3]:
@@
Our strategy for proving Theorem 2.2 will be to show that all the points in $\overline{Nx_{0}}$ are $A$-proximal to $x_{0}$.
Recall our notation for the projection $p:T^{1}\Sigma_{s}\to\Sigma_{s}$. The half planes
$\mathcal{O}=\left\{z\in\mathbb{H}^{2}:\mathrm{Im}z<0\right\}\times\left\{0,1\right\},$
whose boundaries are $p(A_{+}x_{0})$ and $p(A_{+}x_{1})$ respectively, are isometrically embedded in $\Sigma_{s}$. Therefore if a geodesic ray $\ell_{+}=p(A_{+}w)$ in $\Sigma_{s}$ enters $\mathcal{O}$ it can never leave it. We observe the following:
Lemma 2.4 (Bottleneck Lemma).
There exists a sequence $\delta_{k}\to 0$ for which any $w\in\mathsf{T}^{1}\Sigma_{s}$ with $\sup\tau\left(A_{+}w\right)=+\infty$ and $p(A_{+}w)\subseteq\Sigma_{s}\smallsetminus\mathcal{O}$ satisfies
$d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\leq\delta_{k}\quad\text{for all large }k,$
where $t_{k}\to\infty$ is the sequence of times in which $\tau(a_{t_{k}}w)=\tau(a_{s_{k}}x_{0})=s_{k}$.
More generally, if $A_{+}w$ is any geodesic ray with unbounded positive $\tau$-value then
(2.1) $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\to 0,$
for similarly chosen $t_{k}$’s.
This lemma describes a sequence of arbitrarily thin “bottlenecks”, depicted in Figure 3, which every geodesic ray with unbounded positive $\tau$-value in $\Sigma_{s}\smallsetminus\mathcal{O}$ is forced to pass through.
@@
(2.2) $\lim_{k\to\infty}d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{s_{k}}x_{0},a_{s_{k}}x_{1}\right)=0.$
Let $I_{k}=[a_{s_{k}}x_{,}a_{s_{k}}x_{1}]$ be the “vertical” geodesic segment connecting the $a_{s_{k}}x_{i}$’s. Note that $I_{k}=\tau^{-1}(s_{k})\cap\Sigma_{s}\smallsetminus\mathcal{O}$.
Let $w\in\mathsf{T}^{1}\Sigma_{s}$ with $\ell_{+}=p(A_{+}w)\subseteq\Sigma_{s}\smallsetminus\mathcal{O}$ be as in the statement of the lemma. For all large $k$, the balls $B_{k}=B_{3}(a_{s_{k}}x_{0})$ of radius 3 around $a_{s_{k}}x_{0}$ are isometrically embedded in $\Sigma_{s}$ and contain $\mathsf{T}^{1}I_{k}=p^{-1}(I_{k})$. The geodesic segments $A_{[-1,1]}a_{s_{k}}x_{i}$ are contained in $B_{k}$ for all large $k$ and their Hausdorff distance tends to 0 in $\mathsf{T}^{1}\Sigma_{s}$. By our assumption, $\ell_{+}\cap B_{k}\subseteq B_{k}\smallsetminus\mathcal{O}$ implying that $\ell_{+}\cap I_{k}\neq\emptyset$, see Figure 3. One can readily verify that
$\delta_{k}:=\textrm{diam}\left\{y\in\mathsf{T}^{1}I_{k}:p\left(A_{[-1,1]}y\right)\cap p\left(A_{[-1,1]}a_{s_{k}}x_{0}\cup A_{[-1,1]}a_{s_{k}}x_{1}\right)=\emptyset\right\}$
tends to 0 as $k\to\infty$. Since $p\left(A_{+}w\right)\cap p\left(Aa_{s_{k}}x_{0}\cup Aa_{s_{k}}x_{1}\right)=\emptyset$ we have
(2.3) $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{s_{k}}x_{0}\right)\leq\delta_{k}\qquad\text{for all large }k,$
where $a_{t_{k}}w\in\mathsf{T}^{1}I_{k}$, as claimed.
More generally, notice that any geodesic ray $A_{+}w$ with unbounded $\tau$-value is either eventually contained in $\Sigma_{s}\smallsetminus\mathcal{O}$ or is asymptotic to one of the $A_{+}x_{i}$’s. In both cases we conclude (2.1). ∎
As a direct corollary we deduce:
Corollary 2.5.
If $\beta(w)=b>$-$\infty$ then $w$ is $A$-proximal to $a_{b}x_{0}$.
Proof.
Assume $\beta(w)=b>$-$\infty$, then in particular $\tau(a_{t}w)\to+\infty$ as $t\to+\infty$ and hence (2.1) holds for $\tau(a_{t_{k}}w)=\tau(a_{s_{k}}x_{0})$. By the definition of $\beta(w)$, in (1.2), we have
$\lim_{k\to\infty}s_{k}-t_{k}=\lim_{k\to\infty}\tau(a_{t_{k}}w)-t_{k}=\beta(w)=b,$
that is, $d_{\mathsf{T}^{1}\Sigma_{s}}\left(a_{t_{k}}w,a_{t_{k}+b}x_{0}\right)\to 0$, as claimed. ∎
A second useful observation is the following:
@@
For any $\varepsilon>0$ there exists $k_{0}\in\mathbb{N}$ for which all weaving patterns $W$ with $W\geq k_{0}$ satisfy $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon$.
@@
Let $W=\{k_{1},k_{2},$...\} be any weaving pattern with $W\geq k_{0}$. By the additivity and non-negativity of slack we conclude that the “untightened” broken path
$\alpha_{W}=\eta_{k_{1}}^{\pm}\ast\eta_{k_{2}}^{\mp}\ast$...,
given by concatenating the appropriate subsegments of each crossing has
$\mathscr{S}_{+}(\alpha_{W})\leq\sum_{j=k_{0}}^{\infty}\mathscr{S}_{+}(\eta_{j}^{\pm})<\varepsilon.$
The weaving geodesic $\eta_{W,\pm}=\widehat{\alpha_{W}}$ is given by pulling $\alpha_{W}$ tight. By construction, both $\alpha_{W}$ and $\eta_{W,\pm}$ are contained in $\Sigma_{s}\smallsetminus\mathcal{O}$ and have unbounded $\tau$-values, implying by Lemma 2.4 the existence of $p^{+}_{k}\in\eta_{W,\pm}$ and $q^{+}_{k}\in\alpha_{W}$ satisfying
$\tau(p^{+}_{k})=\tau(q^{+}_{k})=s_{k}\to+\infty\qquad\text{and}\qquad d_{\Sigma_{s}}(p^{+}_{k},q^{+}_{k})\to 0.$
Moreover, both paths are backwards asymptotic to one of the $A_{-}x_{i}$’s giving sequences $p^{-}_{k}$ and $q^{-}_{k}$ with equal $\tau$-values tending to -$\infty$ and having $d_{\Sigma_{s}}(p^{-}_{k},q^{-}_{k})\to 0$.
Denote $\eta_{W,\pm}^{k}$ the subsegment of $\eta_{W,\pm}$ connecting $p^{-}_{k}$ to $p^{+}_{k}$, and by $\alpha_{W}^{k}$ the subsegment of $\alpha_{W}$ connecting $q^{-}_{k}$ to $q^{+}_{k}$. By construction, the length of $\eta_{W,\pm}^{k}$ is shorter than the length of $p^{-}_{k},q^{-}_{k}]\ast\alpha_{W}^{k}\ast[p^{+}_{k},q^{+}_{k}$] implying
$\mathscr{S}_{+}(\eta_{W,\pm})$ $=\lim_{k\to\infty}\mathscr{S}_{+}(\eta_{W,\pm}^{k})\leq$
[…]
$=\mathscr{S}_{+}(\alpha_{W})+0+0<\varepsilon.$
We are now set to prove the theorem:
@@
First, by (2.2) we know that $x_{0}$ and $x_{1}$ are $A$-proximal, therefore implying $x_{1}\in F=\overline{Nx_{0}}$ by 2.3. Since $x_{1}\notin Nx_{0}$, as the corresponding $A_{+}x_{j}$’s are not asymptotic, we deduce that $F\neq Nx_{0}$. Recall that horocycles outside of $\mathcal{E}$, those who are based outside the limit set, are properly embedded in the quotient manifold, hence $Nx_{0}\subseteq\mathcal{E}$. Strict inclusion $F\subsetneq\mathcal{E}$ follows, for instance, from $x_{0}$ being quasi-minimizing, see [9, 7].